OpenStudy (anonymous):

A car traveling along Route 66 at a rate pf 75 miles per hour, and the diameter of its wheels are 2.4 feet. Find the number of revolutions per minute the wheels are turning. Round answers to one decimal place.

OpenStudy (radicalsniper):

You should first convert your rate to miles/minute..

OpenStudy (anonymous):

That's correct. OK, will start there.

OpenStudy (radicalsniper):

And then you will use the equation for circumference.

OpenStudy (shane_b):

So you can see the steps... \[\frac{75\cancel{mi}}{hr}*\frac{5280ft}{\cancel{mi}}=\frac{396,000ft}{hr}\]\[\frac{396,000ft}{\cancel{hr}}*\frac{1\cancel{hr}}{60m}=6600ft/m\]\[C_{wheel}=\pi d = \pi (2.4ft)=7.54ft\]\[\frac{6600ft/m}{7.54ft}=\text{~}875.3rpm\]

OpenStudy (anonymous):

Shane, THANKS. Thanks y'all!

OpenStudy (anonymous):

We went over this in class and our professors said we need to us d=2.4 feet and r = 1.2 feet. v=rw w=r/v That's where what y'all today me to convert everything comes in. :)